Please help me solve this equation:
Solve each system addition. Determine whether the equations are independent, dependent, or inconsistent. Solve each equation by the substitution method.
2x = 2 - y
3x + y = -1
Thank you


Answer by mananth(16949)

2x = 2 - y
x= 2/2 -y/2
x=1-y/2
...
plug value of x
3x + y = -1
3(1-y/2)+y=-1
3-3y/2+y=-1
-y/2=-4
-y=-8
y=8
...
2x=2-y
plug value of y
2x=2-8
2x=-6
x=-3
...
only one solution
equations are independant.
..
m.ananth@hotmail.ca